OpenStudy (anonymous):

Need to find an approximate value for x in the equation x^3-e^x=-13/4 Lost on how to go about doing it

OpenStudy (mathmale):

The appropriate choice of approach depends upon what course you're taking. I suspect you're beginning Calculus. Is that correct?

OpenStudy (anonymous):

Actually Differential equations..but I am having a brain freeze on the process for solving this...I can approximate e as being 2.72

OpenStudy (mathmale):

Remember Newton's Method?

OpenStudy (anonymous):

No..but i will look it up.

OpenStudy (mathmale):

"Newton's Method for Approximating Roots."

OpenStudy (mathmale):

Do you remember using any other method for approximating the roots of an equation?

OpenStudy (anonymous):

Calc 1 was 2 years ago...I am an older student (much older) so I my courses are spread out over time...thus I am having o go back and re teach some parts

OpenStudy (mathmale):

I'd like to build on an approach that is familiar to you, if possible. Were I doing the problem myself, I'd definitely use Newton's method.

OpenStudy (mathmale):

shall we go thru Newton's Method for approx. roots? or could you propose an approach more familiar?

OpenStudy (anonymous):

I can try..always up for learning something new or forgotten

OpenStudy (mathmale):

You're given \[x^3-e^x=-\frac{ 13 }{ 4 }\]

OpenStudy (mathmale):

would u pls verify that this is correct?

OpenStudy (anonymous):

The actual equation would be x^3-e^x +13/4 >= 0

OpenStudy (mathmale):

Let's look at this as a function, f(x): \[f(x)=x^3-e^x+\frac{ 13 }{ 4 }\]

OpenStudy (mathmale):

our task is to determine the value or values of x at which this function = 0. Can you agree with that?

OpenStudy (anonymous):

Yes...

OpenStudy (anonymous):

So am looking at 1- f(x) / f(x)' ? For the approximation?

OpenStudy (mathmale):

Please find f '(x) now.

OpenStudy (mathmale):

Yes. Apparently you have done this kind of approx. before.

OpenStudy (anonymous):

3x^2-e^x

OpenStudy (mathmale):

Have you a TI-84 calculator or similar?

OpenStudy (anonymous):

No I was looking back at some notes and saw this..just now.

OpenStudy (anonymous):

Yes

OpenStudy (mathmale):

OK. Please, aritrarily, choose a beginning value (solution value) for x. Interestingly, a random choice often works fine.

OpenStudy (anonymous):

So say x =1?

OpenStudy (mathmale):

Sure.

OpenStudy (anonymous):

OK

OpenStudy (mathmale):

Calculator redy?

OpenStudy (anonymous):

.ok

OpenStudy (mathmale):

Please type in the following: \[1-(x^1-e^1+13/4)/(3(1)^2-e^1)\]

OpenStudy (mathmale):

This follows your formula exactly.

OpenStudy (anonymous):

ok

OpenStudy (mathmale):

Your result?

OpenStudy (anonymous):

-36.38387.....

OpenStudy (mathmale):

I've obtained -4.437. I'll do it again, and encourage you to do the same, until our results are equal. Watch the parentheses carefully as you type in this formula.

OpenStudy (mathmale):

surely is hard to type in the keying instructions. How familiar are you with the TI-84?

OpenStudy (anonymous):

ti -83 plus

OpenStudy (mathmale):